# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v1

@@ Abstract
Image safety classifiers play an important role in identifying and mitigating the spread of unsafe images online (e.g., images including violence, hateful rhetoric, etc.). At the same time, with the advent of text-to-image models and increasing concerns about the safety of AI models, developers are increasingly relying on image safety classifiers to safeguard their models. Yet, the performance of current image safety classifiers remains unknown for real-world and AI-generated images. To bridge this research gap, in this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers. First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.). Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models. Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough in mitigating the multifaceted problem of unsafe images. Also, we find that classifiers trained only on real-world images tend to have degraded performance when applied to AI-generated images. Motivated by these findings, we design and implement a comprehensive image moderation tool called PerspectiveVision, which effectively identifies 11 categories of real-world and AI-generated unsafe images. The best PerspectiveVision model achieves an overall F1-Score of 0.810 on six evaluation datasets, which is comparable with closed-source and expensive state-of-the-art models like GPT-4V. UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI. [^1] Our project website is at https://unsafebench.github.io.
@@ Introduction
Unsafe images that include inappropriate content, such as violence, self-harm, and hate, are prevalent across Web communities like Reddit [20] and 4chan [1]. Their presence poses significant challenges to communities and society at large; they can reinforce stereotypes [29, 37, 70], incite hate and violence [46, 39, 55], and trigger self-harm behaviors [54]. To combat this longstanding problem, online platforms rely on image safety classifiers and human moderators to identify and remove unsafe images from the Web. Image safety classifiers like Q16 [63] and Not-Suitable-For-Work (NSFW) detector [16] are trained on real-world unsafe images and are used for detecting real-world unsafe images that are shared on the Web. For instance, LAION-AI [9], a data provider for machine learning models and datasets scrapped from the Web, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [66] and LAION-5B [65]. Overall, both the Web and publicly available datasets include a non-negligible amount of real-world unsafe images that need to be moderated.
Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [60] and Midjourney [12] are also capable of producing unsafe images. According to a recent study [56], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc. A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [3], which further highlights the risks of these text-to-image models. To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images during the generation process and block generated unsafe images before presenting them to end-users. For example, Stable Diffusion implements a safety filter [21] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images. Additionally, extensive research [62, 38, 76, 31] applies Q16 [63] and NudeNet [17] to evaluate the safety of text-to-image models and propose new safety mechanisms.
Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images. It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance can be generalized to AI-generated images, given that most image safety classifiers are trained on real-world image datasets. Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [48] and GPT-4V [5], provides new possibilities in this field. These models are general-purpose visual language models that have undergone rigorous safety alignment, which empowers them to understand unsafe content portrayed in images. Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
@@
Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a set of models that provide fine-grained classification of unsafe images based on the 11 categories of unsafe images. We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across six external datasets containing both real-world and AI-generated images.
Main Findings. Our study yields the following main findings:
  • Overall, GPT-4V is the top-performing model in identifying unsafe content, with an F1-Score between 0.423 and 0.847 (depending on the category of unsafe content). However, due to its closed-source and commercial nature, widely applying GPT-4V on image safety classification tasks is impractical. Currently, there is a lack of an open-source classifier that can comprehensively and effectively identify the wide spectrum of unsafe images.
@@
  • There is an imbalance in moderation across different types of unsafe images. The Sexual and Shocking categories are more effectively detected, with an average F1-Score close to 0.8. However, the categories of Hate, Harassment, and Self-Harm are not effectively detected, with an average F1-Score below 0.6. For example, certain hateful symbols, such as Neo-Nazi symbols in tattoos, evade detection by GPT-4V.
  • Conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet, experience performance degradation on AI-generated images, with F1-Scores decreasing from 0.650-0.698 to 0.596-0.603. We find that AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout. Through case studies, we verify that these characteristics can indeed affect the classifier’s predictions, thereby contributing to performance degradation on AI-generated images.
  • Leveraging large pre-trained foundation models such as VLMs and CLIP presents higher robustness compared to training a small classifier from scratch. Under perturbed images, VLM-based classifiers present the highest Robust Accuracy (RA) above 0.608. Among conventional classifiers, those that utilize CLIP as an image feature extractor, such as Q16, MultiHeaded, SD_Filter, and NSFW_Detector, demonstrate higher robustness compared to those trained from scratch, e.g., NudeNet. Minimal perturbation is sufficient to deceive NudeNet, with an average RA of only 0.293.
  • Our evaluation shows that PerspectiveVision achieves comparable performance to GPT-4V in identifying unsafe images. The best F1-Score is 0.859 on the UnsafeBench test set and 0.810 on six external evaluation datasets, which include morally negative, nudity, hateful, self-harm, and violent images from either real-world or AI-generated sources. As the top-performing baseline, GPT-4V achieves an F1-Score of 0.818 on these external datasets.
@@
Contributions. Our work makes three important contributions. First, we contribute to the research community with a comprehensive image dataset for image safety research. The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors. It serves as a foundation dataset for future research in this field. Second, we thoroughly assess the effectiveness and robustness of current image safety classifiers and compare them with newly emerged VLMs. Our analysis highlights the varying performance of these classifiers across different unsafe categories, aiding these models to learn the vulnerabilities in identifying specific unsafe content, such as tattooed Neo-Nazi symbols. Furthermore, our analysis shows that AI-generated unsafe images pose challenges to the existing classifiers, especially those trained on real-world images, largely due to the distinct characteristics of AI-generated images. These findings highlight the importance for platform moderators and model developers to include AI-generated images in their training data to enhance models’ generalizability. Finally, inspired by Google’s Perspective API in the NLP domain, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images with comparable performances to GPT-4V. PerspectiveVision is made available as an open-source tool, establishing a baseline for detecting general unsafe images. To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision models. We call for joint and ongoing efforts from the research community to build better content moderation tools.
Ethical Considerations. We curate our dataset based on public databases from LAION-5B and the Lexica website. This process does not involve user deanonymization, therefore, our work is not considered human subjects research by our Institutional Review Boards (IRB). During image annotation, we rely on three authors instead of crowdsourcing workers to avoid exposing unsafe content to third parties. We have also given careful consideration to sharing our dataset responsibly. For example, to prevent misuse, we will make the dataset available on Hugging Face, but only upon request and for research purposes.
@@
Existing Image Safety Classifiers
We introduce five popular image safety classifiers: Q16 [63], MultiHeaded [56], SD_Filter [58], NSFW_Detector [16], and NudeNet [17]. Table 1 provides an overview of their backbone models, training paradigms, and training datasets.
Table 1: Overview of existing image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
@@
Large Visual Language Models (VLMs)
Large visual language models have achieved extraordinary capabilities in understanding visual and text content. Given an image and a text instruction, these models can read the image and generate responses following the instruction. Recent studies [41, 59] show that VLMs can be used to detect user-generated unsafe images [41] and hateful memes [59].
@@
We build UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers. We show an overview of UnsafeBench in Figure 1. Due to the absence of a comprehensive labeled dataset in the image safety domain, we first construct the UnsafeBench dataset, which contains 10K images with human annotations. We then collect the existing image safety classifiers, as conventional classifiers, and three VLMs capable of classifying unsafe images, as VLM-based classifiers. We identify the range of unsafe content covered by these classifiers and align them with our unsafe image taxonomy, i.e., 11 unsafe categories. Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy when faced with perturbed images.
@@
Image Collection. We rely on open large-scale multimodal datasets, LAION-5B [65] and Lexica [10], to collect real-world and AI-generated unsafe images. LAION-5B is currently the largest public image-text dataset in the world [8], containing 5.85 billion image-text pairs collected from webpages [65]. We regard LAION-5B as the source to collect real-world images. [^10] It is possible that webpages also contain AI-generated images. However, LAION-5B was collected in March, 2022 [8], while text-to-image generative models like Stable Diffusion and Midjourney gained popularity after July 2022 [25, 13]. Thus, we believe the majority of the images in LAION-5B are real-world images. Lexica [10] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [60]. Both datasets function as a search engine and enable users to find the most relevant images based on a textual description. Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
@@
To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords. Additionally, we employ a large language model, Vicuna-33b, [^11] https://huggingface.co/lmsys/vicuna-33b-v1.3. to augment the definition of each unsafe category and provide 10 specific examples. We show the generated examples in Table 8 in the Appendix. Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords). We then query LAION-5B and Lexica using these keywords and collect the most relevant images. Specifically, for each query, Lexica returns 50 images by default. LAION-5B, on the other hand, returns a flexible number of image URLs. However, some URLs are repetitive or inaccessible. To ensure a balanced number of images from both datasets, we set the number of returned images from LAION-5B to 100. After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
@@
Image Annotation. We perform a human annotation to determine if these collected images are truly unsafe. Three authors of this paper serve as annotators and manually annotate them independently. We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A). Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts. The annotation process undergoes two rounds. In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image. For images where two annotators disagree, we further introduce the third annotator to provide additional labels. The final label for each image is then determined based on the majority vote among these labels. In the second round, annotators revisit images annotated as unsafe and determine the unsafe category. If an image displays a mix of unsafe elements, we choose the predominant category it violates. Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category.
To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [36, 35], a statistical measure for assessing the agreement among annotators. The score is 0.47, indicating a moderate level of agreement among annotators [36, 35]. We understand such a Fleiss’ Kappa score is not perfect, as the annotators’ perceptions of unsafe images may still vary under the same definition of our unsafe image taxonomy. That explains why it is necessary to introduce the third annotator in the first round to improve the reliability of our annotation.
@@
Dataset Statistics. We show the statistics of our annotated images in Table 2. Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images. For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset.
To the best of our knowledge, the UnsafeBench dataset is currently the most comprehensive annotated dataset in the image safety domain. It consists of 10,146 labeled images, encompassing 11 unsafe categories and two sources (real-world and AI-generated). Each unsafe category contains 769-1,151 images. We display more statistical details in Table 9 and image examples in Figure 10 in the Appendix. In this dataset, each image is equipped with three types of “labels”: 1) if the image is safe or unsafe; 2) if unsafe, it belongs to which unsafe category; and 3) whether it’s from real-world collections or AI-generated content. We believe such characteristics enhance the dataset’s usability. We hope this dataset serves as foundational data and accelerates the development of better open-source image safety classifiers in the era of generative AI.
@@
Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes. To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [49], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain. The uncertain class describes a set of outputs that do not yield a clear prediction. For example, InstructBLIP occasionally describes the content of an image without making a prediction. Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards. We collect several example outputs [^12] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class. for each class and fine-tune RoBERTa in a supervised manner. Ultimately, a VLM-based classifier consists of a VLM and a RoBERTa classifier, which predicts whether an image is safe, unsafe, or uncertain.
@@
Image safety classifiers are designed to target various types of unsafe content. To align the unsafe content covered by a classifier with the taxonomy, i.e., 11 unsafe categories, we first identify the range of unsafe content covered by each classifier by examining its documentation and training data. Then, if any unsafe category from our taxonomy falls into this range, we assign the unsafe category to the classifier. For example, Q16 [63] detects “morally negative” images. We examine the definition of “morally negative” images in [32] and find that it covers a list of unsafe concepts, such as harm, inequality, degradation, etc. By mapping these unsafe concepts to our taxonomy, we align the ambiguous definition of “morally negative” with specific unsafe categories. For VLMs, as we adopt the unsafe image taxonomy from OpenAI, we assume that GPT-4V covers all 11 categories. We also infer that LLaVA and InstructBLIP are capable of handling these categories since they are fine-tuned on the instruction dataset generated by GPT-4V [48, 33]. We show the aligning result in Table 3.
@@
Real-World Images (LAION-5B) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -

### v2

@@ Abstract
With the advent of text-to-image models and concerns about their misuse, developers are increasingly relying on image safety classifiers to moderate their generated unsafe images. Yet, the performance of current image safety classifiers remains unknown for both real-world and AI-generated images. In this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers, with a particular focus on the impact of AI-generated images on their performance. First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.). Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models. Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough to mitigate the multifaceted problem of unsafe images. Also, there exists a distribution shift between real-world and AI-generated images in image qualities, styles, and layouts, leading to degraded effectiveness and robustness. Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which addresses the main drawbacks of existing classifiers with improved effectiveness and robustness, especially on AI-generated images. UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI. [^1] Our webpage is at https://unsafebench.github.io/.
@@ Introduction
Unsafe images that include inappropriate content, such as violence, self-harm, and hate, are prevalent across Web communities like Reddit [18] and 4chan [1]. Their presence poses significant challenges to communities and society at large; they can reinforce stereotypes [26, 34, 67], incite hate and violence [43, 36, 52], and trigger self-harm behaviors [51]. To combat this longstanding problem, online platforms rely on image safety classifiers and human moderators to identify and remove unsafe images from the Web. Image safety classifiers like Q16 [60] and Not-Suitable-For-Work (NSFW) detector [14] are trained on real-world unsafe images and are widely used for detecting unsafe images online. For instance, LAION-AI [8], a data provider for machine learning models, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [63] and LAION-5B [62].
Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [57] and Midjourney [11] are also capable of producing unsafe images. According to a recent study [53], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc. A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [2], which further highlights the risks of these AI products. To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images and block them before presenting them to end-users. For example, Stable Diffusion implements a safety filter [19] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images. Additionally, extensive research [59, 35, 73, 28] applies Q16 [60] and NudeNet [15] to evaluate the safety of text-to-image models and propose new safety mechanisms.
Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images. It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance, i.e., effectiveness and robustness, can be generalized to AI-generated images, given that most are trained on real-world image datasets. Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [45] and GPT-4V [4], provides new possibilities in this field. These models are general-purpose visual language models that have undergone rigorous safety alignment, which empowers them to understand unsafe content portrayed in images. Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
@@
Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [62] dataset for real-world images and the Lexica [9] website for AI-generated images. The UnsafeBench dataset is meticulously annotated and contains 10K images that encompass 11 unsafe categories and two sources (real-world or AI-generated). Next, we collect five common image safety classifiers (Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15]) as conventional classifiers and three VLM-based classifiers built on LLaVA [45], InstructBLIP [30], and GPT-4V [4], combined with a RoBERTa model to classify VLM’s responses. We then examine the specific unsafe content covered by these classifiers and align them with our unsafe image taxonomy. Finally, we assess these classifiers’ effectiveness with the annotated dataset and robustness against random and adversarial perturbations. We particularly focus on the varying performances of classifiers on real-world and AI-generated images. Through clustering techniques and case studies, we identify distinct characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
@@
Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images with enhanced effectiveness and robustness. We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across five external datasets containing both real-world and AI-generated images.
Main Findings. We have the following main findings:
  • Conventional vs. VLM-Based Classifiers. Compared to conventional classifiers, VLMs can identify a wider range of unsafe content, with GPT-4V being the top-performing model. Meanwhile, most conventional classifiers focus only on limited categories of unsafe images. Regarding robustness, although VLM-based classifiers are vulnerable to white-box adversarial attacks, they are comparably more resilient than conventional classifiers. The most vulnerable classifiers are those trained from scratch in a supervised manner without relying on any pre-trained models, like NudeNet.
  • Real-World vs. AI-Generated Images. There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts. This shift leads to the degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet. AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout, which disrupt the classifier’s predictions. Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts. Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
@@
  • Imbalanced Effectiveness. Generally, there is an imbalance in effectiveness across different types of unsafe images. The Sexual and Shocking categories are more effectively detected, with an average F1-Score close to 0.8. However, the effectiveness in the categories of Hate, Harassment, and Self-Harm needs further improvement, with an average F1-Score below 0.6. Although GPT-4V is the top-performing model, it still fails to detect certain hateful symbols, such as Neo-Nazi symbols in tattoos.
  • PerspectiveVision. PerspectiveVision is designed to address the main drawbacks of current image safety classifiers: the most effective model (GPT-4V) being closed-source, imbalanced effectiveness, and degraded performance on AI-generated images. By including a large number of AI-generated images in the training data, PerspectiveVision achieves the highest effectiveness across five external datasets and also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
@@
Contributions. Our work makes three important contributions. First, we contribute to the research community with a comprehensive image dataset for image safety research. The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors. It serves as a foundation dataset for future research in this field. Second, we take the first step of benchmarking the effectiveness and robustness of current image safety classifiers, particularly focusing on the impact of AI-generated images. Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations. These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features like artistic representation and grid layout. Finally, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images across fine-grained categories with enhanced effectiveness and robustness on both real-world and AI-generated images. PerspectiveVision is made available as an open-source tool, establishing a baseline to detect general unsafe images. To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision.
Ethical Considerations. The ethical concern in this study mainly lies in the manual annotation and future release of UnsafeBench images. We have undergone an ethical review by our ethical review board. Our ERB has approved the study and states that there are no ethical considerations if annotators are not exposed to images that are illegal to view or own, such as child sexual abuse materials, which do not exist in our dataset. At the same time, to mitigate any potential harm from the exposure of harmful content, we undertook all annotations ourselves, hence not exposing harmful content to other people. Regarding dataset release, we have also given careful consideration to sharing our dataset responsibly. On the one hand, sharing such a dataset will raise ethical concerns as it contains realistic unsafe images. On the other hand, dataset sharing is important to ensure reproducibility. To account for both ethical concerns and reproducibility, we will make the dataset available but only upon request and for research purposes.
@@
Table 1: Overview of conventional image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
@@
Image Safety Classification Models
Conventional Image Safety Classifiers. Before large models gain popularity, AI practitioners generally rely on smaller, conventional classifiers. These classifiers typically consist of an image feature extractor, such as CLIP, and a head/component that assigns the image to safe/unsafe classes. For example, Q16 [60] is a widely used  [60, 62, 59, 44, 73, 35, 28, 66, 27] binary image classifier that predicts a given image as morally positive or negative. After extracting image features with CLIP, it compares the image feature with two soft prompts (optimized text embeddings), one representing the positive class and the other being the negative class. In this paper, we examine five popular image safety classifiers: Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15]. We summarize the backbone, training paradigm, and training dataset in Table 1 and elaborate on these details in Section A.1 in the Appendix.
Large Visual Language Models. Large visual language models have achieved extraordinary capabilities in understanding visual and text content. Given an image and a text instruction, these models can read the image and generate responses following the instruction. Recent studies [38, 56] show that VLMs can be used to detect user-generated unsafe images [38] and hateful memes [56]. We test both commercial and open-source VLMs including GPT-4V [30], LLaVA (7B) [45], and InstructBLIP (7B) [30]. The specific checkpoints we used are provided in Section A.2 in the Appendix.
@@
We build UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers. We show an overview of UnsafeBench in Figure 1. Due to the absence of a comprehensive labeled dataset in the image safety domain, we first construct the UnsafeBench dataset, which contains 10K images with human annotations. We then collect the existing image safety classifiers, as conventional classifiers, and three VLMs capable of classifying unsafe images, as VLM-based classifiers. We identify the range of unsafe content covered by these classifiers and align them with our unsafe image taxonomy, i.e., 11 unsafe categories. Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy against perturbations.
@@
Image Collection. We rely on open large-scale multimodal datasets, LAION-5B [62] and Lexica [9], to collect real-world and AI-generated unsafe images. LAION-5B is currently the largest public image-text dataset in the world [7], containing 5.85 billion image-text pairs collected from webpages [62]. We regard LAION-5B as the source to collect real-world images. Lexica [9] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [57]. Both datasets serve as a search engine and enable users to find the most relevant images based on a textual description. Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
@@
To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords. Additionally, we employ a large language model, Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3. to augment the definition of each unsafe category and provide 10 specific examples. We show the generated examples in Table 8 in the Appendix. Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords). We then query LAION-5B and Lexica using these keywords and collect the most relevant images. After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
@@
Image Annotation. We perform a human annotation to determine if these collected images are truly unsafe. Three authors of this paper serve as annotators and manually annotate them independently. We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A). Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts. The annotation process undergoes two rounds. In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image. For images where two annotators disagree, we further introduce the third annotator to provide additional labels. The final label for each image is then determined based on the majority vote among these labels. In the second round, annotators revisit images annotated as unsafe and determine the unsafe category. If an image displays a mix of unsafe elements, we choose the predominant category it violates. Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category. To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [33, 32], a statistical measure for assessing the agreement among annotators. The score is 0.47, indicating a moderate level of agreement among annotators [33, 32].
@@
Dataset Statistics. We show the statistics of our annotated images in Table 2 and Table 9. Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images. For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset. To the best of our knowledge, it is currently the most comprehensive annotated dataset in the image safety domain.
Visualizing Distribution Shift. We compare the difference between real-world and AI-generated images in image semantics and qualities, such as signal-to-noise ratio (SNR), edge density, and entropy. Figure 11 and Figure 13 show the different distribution of two groups of images. We find that a distribution shift is present in image semantics, e.g., image embeddings are separated in the 2D space between real-world and AI-generated images. The shift also exists in image qualities, e.g., AI-generated images have lower SNR and higher entropy, indicating they tend to be noisier and more complex.
@@
Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes. To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [46], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain. The uncertain class describes a set of outputs that do not yield a clear prediction. For example, InstructBLIP occasionally describes the content of an image without making a prediction. Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards. We collect several example outputs [^4] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class. for each class and fine-tune RoBERTa in a supervised manner. Ultimately, a VLM-based classifier consists of a VLM and the RoBERTa classifier, which function sequentially.
@@
Image safety classifiers are designed to target various types of unsafe content. To align the unsafe content covered by a classifier with the taxonomy, i.e., 11 unsafe categories, we first identify the range of unsafe content covered by each classifier by examining its documentation and training data. Then, if any unsafe category from our taxonomy falls into this range, we assign the unsafe category to the classifier. For example, Q16 [60] detects “morally negative” images. We examine the definition of “morally negative” content in [29] and find that it covers a list of unsafe concepts, such as harm, inequality, degradation, etc. By mapping these unsafe concepts to our taxonomy, we align the ambiguous definition of “morally negative” with specific unsafe categories. For VLMs, as we adopt the unsafe image taxonomy from OpenAI, we assume that GPT-4V covers all 11 categories. We also infer that LLaVA and InstructBLIP are capable of handling these categories since they are fine-tuned on the instruction dataset generated by GPT-4V [45, 30]. We show the aligning result in Table 3.
Evaluation Methodology
Robustness. We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47]. Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions. The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
We create adversarial examples using three gradient-based adversarial algorithms, FGSM [37], PGD [47], and DeepFool [50], with details introduced in Section A.3 in the Appendix. These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image. For conventional classifiers, we directly use three types of adversarial algorithms to solve the Equation 1 and obtain the optimized perturbations.
However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions. Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task. To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings. For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2. By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label. Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP). Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100. We demonstrate the perturbed images and different types of perturbations in Figure 2.
Figure 2: Perturbed images with Gaussian perturbations and three types of adversarial perturbations bounded by $L_{\infty}$ norm ($\epsilon=0.01$).
Evaluation Metric for Robustness. We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers. RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images, which is also equal to 1 - attack success rate. Here, we generate adversarial examples only for images that are correctly predicted by classifiers during the effectiveness evaluation.
Evaluation Results
@@
LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -